I am lost and confused. Not sure what I am supposed to do.
1)find the x and y intercepts for the graph of the
equation given:
x+7y=0
2) find an equation of the line
containing the given pair of point
(-2,-1) and (-5,-8)
3) find the product of (5z-u)(3z+5u)
4)12 1/2% of what amount is $36.80
5) convert the number to scientific notation
0.0000000008

Answer by rfer(16322)

1)
(0,0)
2)
m+(-8--1)/(-5--2)
m=-7/-3
m=7/3
-------------------
-1=7/3*-2+b
-1=-4.7+b
3.7=b
---------------
y=7/3x+3.7
3)
(5z-u)(3z+5u)
15z^2+25zu-3zu-5u^2
15z^2+22zu-5u^2
4)
36.80/0.125=$294.40
5)
8X10^10